Solve the following system of equations by using the concept of matrices and determinants. πŸ“π’™ + πŸ•π’š + 𝟐 = 𝟎 πŸ’π’™ + πŸ”π’š + πŸ‘ = 𝟎 (b) Find whether the following series are convergent or divergent √ 𝟏 πŸ’ + √ 𝟐 πŸ” + √ πŸ‘ πŸ– + β‹―


1
Expert's answer
2022-03-13T18:45:11-0400

The first item:

"\begin{cases}
 5x + 7y + 2 = 0 \\
 4x + 6y + 3 = 0
 \end{cases}"

"x = \frac{\begin{vmatrix}
 -2 & 7\\
 -3 & 6
 \end{vmatrix}}
 {\begin{vmatrix}
 5 & 7\\
 4 & 6
 \end{vmatrix}} = \frac{-12 + 21}{30 - 28} = \frac{9}2 = 4.5"

"y = \frac{\begin{vmatrix}
 5 & -2\\
 4 & -3
 \end{vmatrix}}
 {\begin{vmatrix}
 5 & 7\\
 4 & 6
 \end{vmatrix}} = \frac{-15 + 8}{30 - 28} = \frac{-7}2 = -3.5"


The second item:

"\frac{\sqrt{1}}{4} + \frac{\sqrt{2}}{6} + \frac{\sqrt{3}}{8} + \dots = \sum_{k=1}^{\infty} \frac{\sqrt{k}}{2k + 2} = \sum_{k=1}^{\infty} \frac{1}{2\sqrt{k} + \frac{2}{\sqrt{k}}} \ge \sum_{k=1}^{\infty} \frac{1}{4\sqrt{k}}"

The last sum series is divergent, so the initial one is divergent too.